Jimmy is towing a basket of sick fish to the veterinarian. The fish and basket have a mass of 109kg, and jimmy is pulling with a force of 1938N at an angle of 37 degrees. How much normal force is acting on the basket?
1
Expert's answer
2019-10-29T11:10:25-0400

We have:


"N+F\sin{37}=mg"

The normal force is acting on the basket:


"N=mg-F\sin{37}=(109)(9.8)-193.8\sin{37}=950\ N"